OpenStudy (anonymous):

Solve the following system of linear equations by using the elimination method. Show all work: 2x = 5y + 4 3x - 2y = -16

OpenStudy (dtan5457):

you need to make both lines have the same value for one of the variables.

OpenStudy (dtan5457):

so first i'd set them both to zero

OpenStudy (dtan5457):

2x-5y-4=0 3x-2y+16=0

OpenStudy (dtan5457):

I'd multiply the first one by -3 the 2nd one by 2

OpenStudy (dtan5457):

(3)2x-5y-4=0 6x-15y-12=0 (-2)3x-2y+16=0 -6x+4y-32=0

OpenStudy (anonymous):

ok then what do i do next

OpenStudy (dtan5457):

Now you add them up

OpenStudy (dtan5457):

6x-15y-12=0 -6x+4y-32=0 Add and you get -11y=44

OpenStudy (dtan5457):

y=?

OpenStudy (anonymous):

-4?

OpenStudy (dtan5457):

yes so now that you know that y=-4 plug that back into the original equation (either one)

OpenStudy (dtan5457):

2x = 5(-4) + 4

OpenStudy (dtan5457):

2x=-20+4 2x=-16 x=?

OpenStudy (anonymous):

-8?

OpenStudy (dtan5457):

exactly so your solution is -8,-4

OpenStudy (anonymous):

awesome thank you!!!

OpenStudy (dtan5457):

Welcome. Feel free to fan and medal :)